# QOS: A Quantum Operating System

### v1

@@ 8.4. Knitter
The Reduce Phase. QOS selects any of the $k$ nodes to perform the reduce step. The rest of the nodes send the intermediate results to this node, which performs a thread-parallel sum of $k$ results. Equivalent to the map phase, the parallel sum can also be executed on GPUs. This produces the final output to be returned to the user (Figure 8 (b), step (2)).
9. Evaluation
9.1. Experimental Methodology
@@
Figure 11. QOS Compiler (§ 9.2). Impact of the QOS compiler on the circuit fidelity against Qiskit (qiskit-transpiler, [n. d.]), CutQC (Tang et al., 2021b), and FrozenQubits (Ayanzadeh et al., 2023). The circuits are optimized using budget $b=3$. There is a mean 2.6×2.6$\times$2.6 ×, 1.6×1.6$\times$1.6 ×, and $1\times1$ improvement for 12-qubit circuits, respectively. There is a 456.5×456.5$\times$456.5 ×, 7.6×7.6$\times$7.6 ×, and 1.67×1.67$\times$1.67 × improvement for circuits of 24 qubits, respectively.
For (1), we use a server with a 64-core AMD EPYC 7713P processor and 512 GB ECC memory. For (2), we conduct our experiments on IBM Falcon r5.11 QPUs. Unless otherwise noted, we use the IBM Kolkata 27-qubit QPU.
Framework and Configuration. We use the Qiskit (Qiskit contributors, 2023) Python SDK for compiling quantum circuits and running simulations. We compile quantum circuits with the highest optimization level (3) and run with 8192 shots. Each data point presented in the figures is the median of five runs.
Benchmarks. We study QOS on a set of circuits used in state-of-the-art NISQ algorithms, adopted from the 3 benchmark suits of Supermarq (Tomesh et al., 2022), MQT-Bench (Quetschlich et al., 2022) and QASM-Bench (Li et al., 2023). The algorithms’ circuits can be scaled by the number of qubits and depth. Specifically. We study 9 benchmarks: GHZ, W-State, Bernstein Vazirani (BV), Hamiltonian Simulation (HS-$t$), Quantum-enhanced Support Vector Machine (QSVM), Two Local Ansatz (TL-$n$), Variational Quantum Eigensolver (VQE-$n$), and Approximate Optimization Algorithm (QAOA-R/P), these benchmarks cover a wide range of relevant criteria for evaluating QOS.
@@
Metrics. We evaluate the following metrics:
  • Fidelity: We use the Hellinger fidelity as a measure of how close a noisy result is to the desired ground truth of a quantum circuit (Hellinger, 1909; fidelity-qiskit, [n. d.]). The Hellinger fidelity is calculated as $Fidelity(P_{ideal},P_{noisy})=\left(1-H\left(P_{ideal},P_{noisy}\right)^{2}\right)^{2}\mapsto[0,1]$, where $H$ is the Hellinger distance between two probability distributions, and $P_{ideal},P_{noisy}$ are the ideal and noisy probability distributions, respectively.
  • Circuit Properties: Number of CNOT gates and depth. When a Qernel contains more than one sub-Qernel, we use the sub-Qernel with the maximum depth, amount of CNOTs, or an average of these two properties.
  • Waiting Time: The time a circuit spends in a QPU’s queue, waiting for execution, in seconds.
  • Classical Overhead: The optimization and post-processing overheads (§ 7) of the QOS compiler vs. Qiskit’s transpiler (qiskit-transpiler, [n. d.]).
  • Quantum Overhead: The number of additional quantum circuits we need to execute per original quantum circuit.
@@
Baselines. We evaluate the QOS compiler against Qiskit v0.41, CutQC (Tang et al., 2021b) and FrozenQubits (Ayanzadeh et al., 2023). QOS’s multi-programmer is evaluated against (Das et al., 2019). Regarding QOS scheduler, to the best of your knowledge, (Ravi et al., 2021b) is the only peer-reviewed quantum scheduler, but it doesn’t provide source code or enough technical details to faithfully implement it.
9.2. QOS Compiler
@@
Figure 12. QOS Compiler (§ 9.2) and QOS Estimator (§ 8.1). (a) Compiler: classical and quantum overheads and fidelity improvement as a relative factor to Qiskit. For 24 qubits, the improvement outweighs the overheads. (b) Compiler: scalability to a large, hypothetical 1000-qubit QPU. Any budget $b>0$ achieves higher quality results than using no optimizations. (c) Estimator’s performance: fidelity of IBM Auckland vs. the QPU automatically selected by the estimator.
@@
RQ1: How well does the QOS compiler improve the fidelity of circuits that run on NISQ QPUs? We evaluate the performance of the QOS compiler w.r.t the post-optimization properties and fidelity of the circuits while also analyzing the classical and quantum costs of our approach.
@@
Effect on the Circuit Depth and Number of CNOTs. In Figure 10, we show the performance of the QOS compiler on the circuits’ depth and number of CNOTs, where we plot the relative difference in post-optimization circuit depth and the number of CNOTs between Qiskit (the red horizontal line) and FrozenQubits (Ayanzadeh et al., 2023), CutQC (Tang et al., 2021b), and the QOS compiler. Figures 10 (a) and (c) show that the circuit depth decreases by $46\%$, $38.6\%$, and $29.4\%$, respectively. Figures 10 (b) and (d) show that the number of CNOTs decreases by $70.5\%$, $66\%$, and $56.6\%$, respectively. The improvement in both metrics against the baselines is attributed to the composability of our compiler; the combined effect of circuit compactions (§ 6) achieves better results than standalone techniques.
@@
Impact on Fidelity. Figure 11 shows the QOS-optimized circuits’ fidelity against Qiskit (qiskit-transpiler, [n. d.]), CutQC (Tang et al., 2021b), and FrozenQubits (Ayanzadeh et al., 2023). The results show a mean 2.6×2.6$\times$2.6 ×, 1.6×1.6$\times$1.6 ×, and $1\times1$ improvement for 12-qubit circuits, respectively, and a 456.5×456.5$\times$456.5 ×, 7.6×7.6$\times$7.6 ×, and 1.67×1.67$\times$1.67 × improvement for circuits of 24 qubits, respectively. The fidelity improvement is a consequence of lower circuit depths and fewer CNOTs, as shown in Figure 10.
@@
Classical and Quantum Overheads. Figure 12 (a) shows the average classical and quantum overheads of the QOS compiler. The classical overhead is 16.6×16.6$\times$16.6 × and 2.5×2.5$\times$2.5 × for 12 and 24 qubits, respectively, and the quantum overhead is 31.3$\times$ and 12$\times$ for 12 and 24 qubits, respectively. However, fidelity improves by 2.6×2.6$\times$2.6 × and 456.5×456.5$\times$456.5 × for 12 and 24 qubits, respectively; therefore, for larger circuits, the fidelity improvement is worth the cost.
@@
Scalability. To demonstrate that the QOS Compiler increases the scalability, we run the VQE-1 benchmark on a hypothetical 1000-qubit QPU with one-qubit gate errors of $10^{-4}$, two-qubit gate errors of $10^{-3}$, and measurement errors of $10^{-2}$. We optimize with budget $b\in\{0,1,4,8\}$ and report the estimated fidelity. Figure 12 (b) shows that all budget $b$ values improve the estimated fidelity, with a tradeoff of improvement vs. overheads.
@@
RQ1 takeaway: The QOS compiler improves the properties of quantum circuits by $51\%$ on average, leading to an improvement in fidelity of 2.6–456.5$\times$, while incurring acceptable classical and quantum overheads.
9.3. Estimator

### v2

@@ 5.2 Post-Processor
The reduce phase. QOS selects any of the $k$ nodes to perform the reduce step. The rest of the nodes send the intermediate results to this node, which performs a thread-parallel sum of $k$ results. Equivalent to the map phase, the parallel sum can also be executed on GPUs. This produces the final output to be returned to the user.
6 Estimator
The estimator is responsible for predicting the fidelity performance of a given Qernel on the underlying QPUs without executing the Qernel, which would be extremely costly. This prediction will be the leading decision factor for the scheduler when assigning the Qernel to a QPU.
Challenges. Estimating fidelity at-priori faces three key challenges: (1) Fidelity is a non-deterministic metric affected by the hardware’s probabilities of errors, which change across QPUs and time unpredictably (§ 3). (2) Simulating a Qernel to obtain fidelity estimates is intractable since the complexity of simulating quantum systems scales exponentially with the number of qubits [22]. (3) Fidelity can be approximated numerically [47]; however, there is an accuracy-cost tradeoff. The more complex the analytical model becomes, the higher the runtime overhead of estimation.
Key idea. The estimator leverages analytical and regression models that do not require real execution or simulation of the Qernel. Specifically, the estimator’s models compute a score for each Qernel-QPU assignment that captures the potential fidelity of that assignment, and the estimator supports configurable scoring policies with different runtime-cost tradeoffs. These policies consider (1) the Qernels’ properties generated from the error mitigator (2) the QPUs’ calibration data (i.e., noise model), which are available to quantum cloud providers since they perform the calibration cycles.
For (1), important properties include the number and types of gates, depth, and the number of measurements (§ 1). For (2), recall that QPUs are characterized by calibration data that describe the exact error rates of the QPU for that calibration cycle (§ 2.2), specifically, the individual qubit readout errors, the individual gate errors, and the $T2$ coherence times. In this work, we implement two scoring policies: a numerical approach for fine-grained control over the estimations and a regression model approach for abstracting away the complexity of estimation.
Numerical cost policy. This policy estimates execution fidelity by transpiling the circuit for the target QPU (§ 2.2). Target transpilation enables fine-grained fidelity estimation by producing the mapping between logical and physical qubits and the gate (instruction) schedule. The mapping captures the expected readout and gate errors, while the gate schedule captures the order and exact timing that the gates will be applied on the qubits, which reveals the hardware decoherence and crosstalk errors, as explained in § 2.2. This method has been explored before, and there are multiple variants, but we base our implementation on Mapomatic [47].
To compute the score, we need the following information: (1) The readout/measurement error probabilities that describe the probability of a bit-flip during measurement, (2) the gate error probabilities, (3) the T2 time, which measures for how long a qubit can stay in an excited state (i.e., |$\ket{1}$), thus quantifying the decoherence error probability, and (4) the crosstalk error which is measured through crosstalk characterization [45]. With this information, we can multiply the probabilities of the respective sources of errors to get an estimate of the total error/fidelity.
Formally, for each qubit $q_{i}$ the readout error is $e_{r(i)}$, for each gate $g_{j}$ the error is $e_{g(j)}$, and the decoherence error is $e_{d(t)}=1-e^{-t/T2_{i}}$, where $t$ is the idle time of the qubit $q_{i}$ (no gates act on it [14]) and $T2$ is the decoherence time of $q_{i}$. The crosstalk error between gates $g_{k}$ and $g_{l}$ is $e_{ct(k,l)}$. Putting it all together, the final fidelity score is computed as follows: $fid=1-\prod_{i=0}^{N}e_{r(i)}e_{d(i)}\prod_{j=0}^{M}e_{g(j)}\prod_{j=0,k=0}^{M\times M}e_{ct(j,k)}$, where $N$ is the circuit’s number of qubits and $M$ is the number of gates. Since all hardware error information is known at-priori, and quantum errors accumulate multiplicatively, this policy produces high-accuracy estimations, as we show in § 9.3.
Figure 7: QOS multi-programmer example workflow (§ 7). (a) We use the estimator’s output to find Qernels with the same best QPU. (b) We compute their independent utilization, and (c) their compatibility score. If compatible, (d) we check for layout overlap, and (e) apply the appropriate multi-programming policy.
Regression model policy. Since the impact of noise errors on fidelity during quantum computation can be described numerically, we can train a regression model to predict the fidelity of a transpiled Qernel on a possible QPU using the QPU’s calibration data and the Qernel’s static properties as features. Specifically, we use the aforementioned errors we defined in the numerical cost policy as QPU features and the static properties (§ 1) as Qernel features. Even simple regression models such as linear regression achieve high prediction accuracy, up to 99%. Although this policy is simple to use without detailed knowledge of the relationship between errors, in QOS, we use the numerical cost policy by default for estimation to have a clear understanding and full control of the process.
7 Multi-programmer
The size of quantum programs that run with high fidelity is small, leading to QPU underutilization (§ 3.3). To increase QPU utilization, QOS multi-programs two or more Qernels, potentially from different users, to run on the same QPU. We refer to this multi-programming as bundling the Qernels together.
Challenges. Bundling Qernels faces three distinct challenges: (1) Trivially bundling Qernels together will deteriorate fidelity because qubits interfere with each other via crosstalk errors (§ 2.2). (2) On top of that, bundled Qernels with unequal runtimes do not optimally increase utilization since effective utilization is measured in space and time (number of QPU qubits used and duration that they are non-idle). (3) Qernel compilation involves several NP-hard processes [70], so we need to minimize it.
Key ideas. To tackle these challenges, we introduce new ideas unexplored by prior multi-programming work [16, 40]. First, we define a new utilization metric, namely effective utilization, that captures the temporal dimension of utilization, i.e., H/W qubit usage during computation time. Next, we compute Qernel compatibility functions that quantify how well-suited any two Qernels are to run together, to minimize fidelity loss and re-compilation times. Last, we create buffer zones between two bundled Qernels to minimize crosstalk errors.
7.1 Qernel Compatibility
Qernel compatibility quantifies crosstalk errors and the effective utilization of bundled Qernels by considering the Qernels’ static properties (§ 1).
Effective utilization. The trivial way to compute utilization in the case of bundled Qernels is by diving their total number of qubits over the QPU’s number of qubits, $N_{total}/N_{QPU}$. However, this is not accurate when the bundled Qernels have unequal runtimes. In order to accurately quantify QPU usage in the context of multi-programming, we define effective utilization as spatial plus temporal utilization. To quantify spatial utilization, it suffices to compute the ratio of allocated QPU qubits (of the longest Qernel in terms of depth) over the QPU’s number of qubits.
To quantify temporal utilization, instead of simply adding the spatial utilization of each allocated Qernel to the total utilization, we first scale it with a weight that is the ratio of the Qernel’s duration over the longest Qernel’s duration. Recall that, to measure the Qernels’ durations, we use the depth property that reflects the longest sequence of gates the Qernel consists of (§ 2.2).
More technically, we define effective utilization as $u_{eff}=\frac{N_{C_{max}}}{N_{QPU}}*100+\sum_{k=1}^{n}\frac{D_{k}}{D_{max}}*\frac{N_{C_{k}}}{N_{QPU}}*100$, where $N_{C_{max}},N_{QPU}$ are the number of qubits of the longest Qernel and the QPU, respectively, $k$ is the number bundled Qernels excluding the longest Qernel, and $D$ is the depth of the Qernel. The first term, $\frac{N_{C_{max}}}{N_{QPU}}*100$, captures the spatial utilization. The right term (sum) captures the temporal dimension, and $\frac{D_{n}}{D_{max}}$ is the relative depth of Qernel $k$ compared to the longest Qernel, and $\frac{N_{C_{k}}}{N_{QPU}}$ is the fraction of QPU qubits used by Qernel $k$.
For example, a 10-qubit Qernel $Q_{0}$ spatially utilizes 50% of a 20-qubit QPU. Now assume that $Q_{0}$ runs 3$\times$ longer than a 10-qubit Qernel $Q_{1}$. During $\frac{2}{3}$ of $Q_{0}$’s runtime, the qubits allocated to $Q_{1}$ will be idle, decreasing the effective utilization to only 66%.
Quantifying crosstalk. To quantify crosstalk without running the bundled Qernels, we use the entanglement ratio and parallelism Qernel feature vectors of [76], where higher values indicate a higher chance for crosstalk errors (§ 2.2). Intuitively, the entanglement ratio captures the proportion of 2-qubit gates over all gates, and parallelism captures how many gates run in parallel per time unit. Thus, more parallel 2-qubit gates equal a higher chance of crosstalk errors [45].
Compatibility formula. To quantify the compatibility of two candidate Qernels, we consider their effective utilization, as well as their joint entanglement ratio and parallelism scores. This is because we ideally want high effective utilization (not just higher spatial) and fewer crosstalk-induced errors.
Formally, we score a possible Qernel pair as follows: $qc=\alpha~{}u_{eff}+\beta~{}ER_{b}+\gamma~{}PA_{b}\mapsto[0,1]$. Higher score is better, $\alpha+\beta+\gamma=1$ (the weights add up to one), $u_{eff}$ is effective utilization, $ER$ is entanglement ratio, $PA$ is parallelism, and $b$ denotes bundled, i.e., $E_{b}$ is the entanglement ratio of the bundled Qernels. The four variables are tunable to give priorities on different objectives, e.g., prioritize effective utilization or minimize crosstalk. After experimenting and fine-tuning, we found that $\alpha=0.25,\beta=0.25,\gamma=0.5$, and $qc\geq 0.75$ gives balanced results, as we show in § 9.4.
Buffer zone. Crosstalk characterization studies on real machines have shown that the probability of crosstalk errors drops exponentially with the distance between any two qubits [45]. We leverage this to minimize crosstalk errors by creating a buffer zone between any two bundled Qernels, i.e., at least one hardware qubit between the two sets of allocated qubits must be free. Given the hardware constraints and limited number of available qubits, we limit the buffer size to up to two qubits of distance.
7.2 Multi-programming Policies
In this work, we implement two multi-programming policies; the first is the fast path multi-programming while the second requires re-compilation and re-estimation.
Restrict policy. The restrict policy checks if there is no overlap in the Qernel mappings on the QPU, including the buffer zone. For example, in Figure 7 (d) in the left case, the logical qubits of the Qernels are mapped to disjoint sets of physical qubits on the QPU, and there is a buffer zone between the two sets. In that case, the policy simply bundles the Qernels together, and fidelity loss is minimized through the aforementioned compatibility score.
Re-evaluation policy. This policy is the fallback of the restrict policy. If the Qernel mappings overlap, the two Qernels are transpiled again for the target QPU, and their new fidelity is estimated. If the new fidelity is lower up to a fixed $\epsilon>0$ value compared to the original independent fidelities, the bundling is kept. Otherwise, the multi-programmer selects the next most compatible Qernel pair.
Example. Figure 7 shows an example where the multi-programmer receives three Qernels with three estimations each and identifies $Qernel~{}0$ and $Qernel~{}2$ as a possible pair since their best QPU is the same ($QPU5$) (a). It computes their independent utilization (31% and 37%, respectively), and the combined utilization is under 100% (b). It computes the compatibility score that surpasses the threshold ($0.9>0.75$) (c). In (d), the right case shows an overlap in the buffer zone (the red qubit), so we apply the re-evaluation policy.
Unbundling. To unbundle the results, the multi-programmer keeps a record that maps the initial (solo) Qernel IDs to the new, bundled Qernel ID, as well as the Qernels’ sizes. Therefore, when receiving a new result from a Qernel with an ID $i$, it scans the record to find an entry $i$, and if found, it splits the probability distribution bitstrings (§ 2.1) into two parts: the left-most and the right-most bits based on the Qernel sizes.
8 Scheduler
Figure 8: Error mitigator (§ 9.2). Impact of the error mitigator on the circuit depth and the number of CNOTs. The circuits are optimized using budget $b=3$, and we compare against Qiskit (red horizontal line), FrozenQubits [4] and CutQC [73]. There is an average $46\%$, $38.6\%$, and $29.4\%$ reduction in circuit depth, and $70.5\%$, $66\%$ and $56.6\%$ reduction in the number of CNOTs, respectively.
Scheduling quantum programs involves fundamental tradeoffs between conflicting objectives; specifically, users ideally want the fidelity of the best fidelity QPU and the waiting time of least busy QPU.
Challenges. Scheduling quantum tasks faces three key challenges: (1) To maximize fidelity, most programs must be scheduled on the same subset of best-performing QPUs (§ 3.2), forming hotspots and necessarily increasing user waiting times. (2) The scheduler must at least know estimates of the quantum task durations. However, similar to fidelity estimation (§ 6), quantum execution time estimation must be fast, without real execution or simulation. (3) Job scheduling is a well-known NP-hard problem; therefore, every heuristic or greedy solution will present a tradeoff between optimality and performance.
Key ideas. Our scheduler leverages an analytical model to estimate the Qernel execution time quickly. Then, based on the fidelity predictions from the estimator, it assigns and runs Qernels across space (which QPUs) and time (when). Last, our scheduler supports configurable policies for managing the aforementioned tradeoffs, prioritizing maximal fidelity, minimal waiting times, or a balanced approach. Our two main policies are a fast heuristic algorithm and a genetic multi-objective optimization algorithm.
Execution time estimation. To optimize for minimal waiting times, the scheduler must first estimate each Qernel’s execution time and then aggregate the execution time estimations in each QPU’s queue to compute the total waiting times. To estimate the execution time, we iterate the longest path (depth) of a Qernel (§ 1) that corresponds to the longest-duration gate chain and thus defines the Qernel’s execution time. By summing the gate durations of each node in the longest path, we get the Qernel’s total execution time.
Formula-based policy. Optimizing for conflicting objectives involves comparing two possible solutions (e.g., maximal fidelity vs. minimal waiting times). In the formula-based policy, we use the following scoring formula: Score = $c\frac{f_{2}-f_{1}}{f_{1}}-(1-c)\frac{t_{2}-t_{1}}{t_{1}}+\beta\frac{u_{2}-u_{2}}{u_{1}}$ to compare and select between two possible assignments. This formula factors fidelity, waiting time, and utilization to determine which assignment is better, given priorities. The parameters are as follows: $f_{i}$: fidelity of the estimation result $i$, $t_{i}$: waiting time for the QPU from estimation result $i$, $u_{i}$: utilization of the QPU for estimation result $i$, $c\in(0,1)$: a system-defined constant that weighs the fidelity difference between estimations and finally, $\beta$: a system-defined constant acting as a weighting factor for utilization difference, balancing system throughput and fidelity. By selecting higher $c$, the system prioritizes fidelity over waiting times, and vice versa, and by selecting higher $\beta$ the system prioritizes utilization over fidelity, and vice versa. By default, $c=\beta=0.5$, which aims for balanced fidelity, waiting times, and utilization.
Genetic algorithm policy. Genetic algorithms excel at optimizing for conflicting objectives by efficiently searching over vast search spaces, and for that, they can be used in the context of QOS. We formulate a multi-objective optimization problem with the conflicting objectives of fidelity vs. waiting times and use the NSGA-II genetic algorithm [17] to solve it. The algorithm creates a Pareto front of possible solutions (schedules), each achieving a different combination of average fidelity and average waiting times. Then, to select one of those schedules, we use the aforementioned formula to score each schedule and select the schedule with the highest score.
9 Evaluation
9.1 Experimental Methodology
@@
Figure 9: Error mitigator (§ 9.2). Impact of the error mitigator on the circuit fidelity against Qiskit [58], CutQC [73], and FrozenQubits [4]. The circuits are optimized using budget $b=3$. There is a mean 2.6×2.6$\times$2.6 ×, 1.6×1.6$\times$1.6 ×, and $1\times1$ improvement for 12-qubit circuits, respectively. There is a 456.5×456.5$\times$456.5 ×, 7.6×7.6$\times$7.6 ×, and 1.67×1.67$\times$1.67 × improvement for circuits of 24 qubits, respectively.
For (1), we use a server with a 64-core AMD EPYC 7713P processor and 512 GB ECC memory. For (2), we conduct our experiments on IBM Falcon r5.11 QPUs. Unless otherwise noted, we use the real Kolkata 27-qubit QPU hosted by IBM.
Framework and configuration. We use the Qiskit [33] Python SDK for compiling quantum circuits and running simulations. We compile quantum circuits with the highest optimization level (3) and run with 8192 shots. Each data point presented in the figures is the median of five runs.
Benchmarks. We study QOS on a set of circuits used in state-of-the-art NISQ algorithms, adopted from the 3 benchmark suits of Supermarq [76], MQT-Bench [60] and QASM-Bench [39]. The algorithms’ circuits can be scaled by the number of qubits and depth. Specifically. We study 9 benchmarks: GHZ, W-State, Bernstein Vazirani (BV), Hamiltonian Simulation (HS-$t$), Quantum-enhanced Support Vector Machine (QSVM), Two Local Ansatz (TL-$n$), Variational Quantum Eigensolver (VQE-$n$), and Approximate Optimization Algorithm (QAOA-R/P), these benchmarks cover a wide range of relevant criteria for evaluating QOS.
@@
Metrics. We evaluate the following metrics: (1) Hellinger Fidelity. As defined in [21], where it ranges in $[0,1]$ and higher is better. (2) Static Properties. Number of CNOT gates and depth (§ 2.2). (3) Waiting Time. The time a Qernel spends in a QPU’s queue, waiting for execution, in seconds. (4) Classical Overhead. The error mitigation classical overheads as a factor of runtime increase ($\times$). (5) Quantum Overhead. The error mitigation quantum overheads as a factor of numbers of circuits increase ($\times$).
@@
Baselines. We evaluate the error mitigator against the standalone fidelity-improving frameworks Qiskit v0.41, CutQC [73] and FrozenQubits [4] to show the efficacy of the error mitigator’s composability. We evaluate QOS’s multi-programmer against [16], the state-of-the-art multi-programming framework. Regarding QOS scheduler, to the best of your knowledge, [62] is the only peer-reviewed quantum scheduler, but it doesn’t provide source code or enough technical details to faithfully implement it. Lastly, we do not perform end-to-end evaluation against a unified baseline since simply combining prior works into a pipeline does not capture the cross- and within-layer optimizations that QOS performs.
9.2 QOS Error Mitigator
@@
Figure 10: Error mitigator (§ 9.2) and estimator (§9.3). (a) Error mitigator: classical and quantum overheads and fidelity improvement as a relative factor to Qiskit. For 24 qubits, the improvement outweighs the overheads. (b) Error mitigator: scalability to a large, hypothetical 1000-qubit QPU. Any budget $b>0$ achieves higher quality results than using no optimizations. (c) Estimator’s performance: fidelity of IBM Auckland vs. the QPU automatically selected by the estimator.
@@
RQ1: How well does the error mitigator improve the fidelity of circuits that run on NISQ QPUs? We evaluate the performance of the error mitigator w.r.t the post-mitigation properties and fidelity of the circuits while also analyzing the classical and quantum costs of our approach.
@@
Effect on the circuit depth and number of CNOTs. In Figure 8, we show the performance of the error mitigator on the circuits’ depth and number of CNOTs, where we plot the relative difference in post-optimization circuit depth and the number of CNOTs between Qiskit (the red horizontal line) and FrozenQubits [4], CutQC [73], and our error mitigator. Figures 8 (a) and (c) show that the circuit depth decreases by $46\%$, $38.6\%$, and $29.4\%$, respectively. Figures 8 (b) and (d) show that the number of CNOTs decreases by $70.5\%$, $66\%$, and $56.6\%$, respectively. The improvement in both metrics against the baselines is attributed to the composability of our error mitigator, where the combined techniques achieve better results than any standalone technique. In cases where QOS under-performs compared to the baselines (e.g., BV benchmark for 24 qubits), it is because the error mitigator achieved the desired circuit size with the qubit reuse technique, which incurs additional costs (§5.1).
@@
Impact on fidelity. Figure 9 shows the error-mitigated circuits’ fidelity against Qiskit [58], CutQC [73], and FrozenQubits [4]. The results show a mean 2.6×2.6$\times$2.6 ×, 1.6×1.6$\times$1.6 ×, and $1\times1$ improvement for 12-qubit circuits, respectively, and a 456.5×456.5$\times$456.5 ×, 7.6×7.6$\times$7.6 ×, and 1.67×1.67$\times$1.67 × improvement for circuits of 24 qubits, respectively. The fidelity improvement is a consequence of lower circuit depths and fewer CNOTs, as shown in Figure 8.
@@
Classical and quantum overheads. Figure 10 (a) shows the average classical and quantum overheads of the error mitigator. The classical overhead is 16.6×16.6$\times$16.6 × and 2.5×2.5$\times$2.5 × for 12 and 24 qubits, respectively, and the quantum overhead is 31.3$\times$ and 12$\times$ for 12 and 24 qubits, respectively. However, fidelity improves by 2.6×2.6$\times$2.6 × and 456.5×456.5$\times$456.5 × for 12 and 24 qubits, respectively; therefore, for larger circuits, the fidelity improvement is worth the cost.
@@
Scalability. To demonstrate that the error mitigator is scalable, we run the VQE-1 benchmark on a hypothetical 1000-qubit QPU with one-qubit gate errors of $10^{-4}$, two-qubit gate errors of $10^{-3}$, and measurement errors of $10^{-2}$. We optimize with budget $b\in\{0,1,4,8\}$ and report the estimated fidelity. Figure 10 (b) shows that all budget $b$ values improve the estimated fidelity, where higher $b$ equals higher fidelity.
@@
RQ1 takeaway. The error mitigator improves the properties of quantum circuits by $51\%$ on average, leading to an improvement in fidelity of 2.6–456.5$\times$, while incurring justifiable classical and quantum overheads.
9.3 Estimator